# Can Watermarks Survive Translation? On the Cross-lingual Consistency of Text Watermark for Large Language Models

### v1

@@ Abstract
Text watermarking technology aims to tag and identify content produced by large language models (LLMs) to prevent misuse. In this study, we introduce the concept of “cross-lingual consistency” in text watermarking, which assesses the ability of text watermarks to maintain their effectiveness after being translated into other languages. Preliminary empirical results from two LLMs and three watermarking methods reveal that current text watermarking technologies lack consistency when texts are translated into various languages. Based on this observation, we propose a Cross-lingual Watermark Removal Attack (CWRA) to bypass watermarking by first obtaining a response from an LLM in a pivot language, which is then translated into the target language. CWRA can effectively remove watermarks by reducing the Area Under the Curve (AUC) from 0.95 to 0.67 without performance loss. Furthermore, we analyze two key factors that contribute to the cross-lingual consistency in text watermarking and propose a defense method that increases the AUC from 0.67 to 0.88 under CWRA.
@@ 1 Introduction
Large language models (LLMs) like GPT-4 OpenAI (2023) have demonstrated remarkable content generation capabilities, producing texts that are hard to distinguish from human-written ones. This progress has led to concerns regarding the misuse of LLMs, such as the risks of generating misleading information, impersonating individuals, and compromising academic integrity Chen and Shu (2023a, b). As a countermeasure, text watermarking technology for LLMs has been developed, aiming at tagging and identifying the content produced by LLMs Kirchenbauer et al. (2023a); Liu et al. (2023b, c). Generally, a text watermarking algorithm embeds a message within LLM-generated content that is imperceptible to human readers, but can be detected algorithmically. By tracking and detecting text watermarks, it becomes possible to mitigate the abuse of LLMs by tracing the origin of texts and ascertaining their authenticity.
@@
The robustness of watermarking algorithms, i.e., the ability to detect watermarked text even after it has been modified, is important. Recent works have shown strong robustness under text rewriting and copy-paste attacks Liu et al. (2023b); Yang et al. (2023). However, these watermarking techniques have been tested solely within monolingual contexts. In practical scenarios, watermarked texts might be translated, raising questions about the efficacy of text watermarks across languages (see Figure 1). For example, a malicious user could use a watermarked LLM to produce fake news in English and then translate it into Chinese. Obviously, the deceptive impact persists regardless of the language, but it is uncertain whether the watermark would still be detectable after such a translation. To explore this question, we introduce the concept of cross-lingual consistency in text watermarking, aiming to characterize the ability of text watermarks to preserve their strength across languages. Our preliminary results on 2 LLMs $\times$ 3 watermarks reveal that current text watermarking technologies lack consistency across languages.
@@
In light of this finding, we propose the Cross-lingual Watermark Removal Attack (CWRA) to highlight the practical implications arising from deficient cross-lingual consistency. When performing CWRA, the attacker begins by translating the original language prompt into a pivot language, which is fed to the LLM to generate a response in the pivot language. Finally, the response is translated back into the original language. In this way, the attacker obtains the response in the original language and bypasses the watermark with the second translation step. CWRA outperforms re-writing attacks, such as re-translation and paraphrasing Liu et al. (2023c), as it achieves the lowest AUC (reducing it from 0.95 to 0.67) and the highest text quality.
@@
To resist CWRA, we propose a defense method that improves the cross-lingual consistency of current LLM watermarking. Our method is based on two critical factors. The first is the cross-lingual semantic clustering of the vocabulary. Instead of treating each token in the vocabulary as the smallest unit when ironing watermarks, as done by KGW Kirchenbauer et al. (2023a), our method considers a cluster of tokens that share the same semantics across different languages as the smallest unit of processing. In this way, the post-translated token will still carry the watermark as it would fall in the same cluster as before translation. The second is cross-lingual semantic robust vocabulary partition. Inspired by Liu et al. (2023b), we ensure that the partition of the vocabulary are similar for semantically similar contexts in different languages. Despite its limitations, our approach elevates the AUC from 0.67 to 0.88 under the CWRA, paving the way for future research.
@@
  • We reveal the deficiency of current text watermarking technologies in maintaining cross-lingual consistency (§ 3).
  • Based on this finding, we propose CWRA that successfully bypasses watermarks without degrading the text quality (§ 4).
  • Based on our analysis of the two key factors for improving cross-lingual consistency, we propose a defense method against CWRA (§ 5).
@@ 2.2 Watermarking for LMs
$s=(|\boldsymbol{x}|_{g}-\gamma|\mathcal{V}|)/\sqrt{|\mathcal{V}|\gamma(1-\gamma)},$ (2)
@@ 3.1 Definition
We define cross-lingual consistency as the ability of a watermark, embedded in a text produced by an LLM, to retain its strength after the text is translated into another language. We represent the original strength of the watermark as a random variable, denoted by $S$ (§ A.1), and its strength after translation as $\hat{S}$. To quantitatively assess this consistency, we employ the following two metrics.
@@
A lower RE indicates that the watermark retains strength close to its original value after translation, signifying great cross-lingual consistency. To avoid the instability caused by values of $S$ that are close to 0, we first aggregate the data by text length and replace the original values of $S$ and $\hat{S}$ with their respective mean values within each group. We also apply min-max normalization to ensure that all values are non-negative.
@@ 3.2 Experimental Setup
We sampled a subset of 1,000 prompts from the UltraChat test set Ding et al. (2023) [^2] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k, and generated responses from the LLM using the text watermarking methods described in § 2.2. The default decoding method was multinomial sampling, and both the prompts and the LLM-generated responses were in English. To evaluate the cross-lingual consistency, these watermarked responses were translated into four languages using gpt-3.5-turbo-0613333https://platform.openai.com/docs/models: Chinese (Zh), Japanese (Ja), French (Fr), and German (De). Notably, English shares greater similarities with French and German, in contrast to its significant differences from Chinese and Japanese.
@@
  • Baichuan-7B Baichuan. (2023): an LLM trained on 1.2 trillion tokens. It offers bilingual support for both Chinese and English.
  • Llama-2-7b-chat Touvron et al. (2023): trained on 2 trillion tokens and only provides support for English.
Method PCC normal-$\uparrow$ RE (%) normal-$\downarrow$
@@
Figure 2: Trends of watermark strengths with text length before and after translation. This is the average result of Baichuan-7B and Llama-2-7b-chat. Figure 7 displays results for each model. Given the distinct calculations for watermark strengths of the three methods, the y-axis scales vary accordingly.
@@ 3.3 Results
Table 1 presents the results, and Figure 2 illustrates the trend of watermark strengths with text length.
@@
Only SIR exhibits such a characteristic: when using Llama-2-7b-chat, its cross-lingual consistency performs notably better among similar languages compared to distant ones, especially in terms of PCC. This can be attributed to its semantic invariance and shared words among similar languages. However, this characteristic is not shown on Baichuan-7B, which might be related to tokenization. In contrast, this property is less evident in the case of KGW and UW.
@@
Overall, SIR indeed exhibits superior cross-lingual consistency compared to KGW and UW. Particularly when using Baichuan-7B, SIR achieves the best PCCs for all target languages. When using Llama-2-7B-chat, SIR still performs well in languages similar to English. This finding highlights the importance of semantic invariance in preserving watermark strength across languages, which we will explore more in § 5. Despite its superiority, SIR still presents a notable reduction in watermark strength in cross-lingual scenarios, as evidenced by Figure 2e.
@@ 4.2 Results
CWRA demonstrates the most effective attack performance, significantly diminishing the AUC and the TPR. For one thing, existing watermarking techniques are not designed for cross-lingual contexts, leading to weak cross-lingual consistency. For another thing, strategies such as Re-translation and Paraphrase are essentially semantic-preserving text rewriting. Such strategies tend to preserve some n-grams from the original response, which may still be identifiable by the watermark detection algorithm. In contrast, CWRA reduces such n-grams due to language switching.
@@
Figure 4: ROC curves for KGW, UW, and SIR under various attack methods: Re-translation, Paraphrase and CWRA. We also present TPR values at a fixed FPR of 0.1. This is the overall result of text summarization and question answering. Figure 8 and Figure 9 display results for each task.
@@ 5.1 Two Key Factors of Cross-lingual Consistency
KGW-based watermarking methods fundamentally depend on the partition of the vocabulary, i.e., the red and green lists, as discussed in § 2.2. Therefore, cross-lingual consistency aims to ensure that the green tokens in the watermarked text will still be recognized as green tokens after being translated into other languages.
With this goal in mind, we start our analysis with a simple English-Chinese case in Figure 5:
@@
    Given “I watch” as the prefix, a watermarked LM predicts “movies” as the next token. Due to watermarking, “movies” is a green token. All candidate tokens are [movies, birds, 电影, 鸟], where “movies” and “电影” are semantically equivalent, as are “birds” and “鸟”.
@@
    A machine translator (MT) then translates the entire sentence “I watch movies” into Chinese: “我 看 电影”.
The question of interest is: what conditions must the vocabulary partition meets so that the token “电影”, the Chinese equivalent of “movies”, also falls within the green list?
@@
    Cross-lingual semantic clustering of the vocabulary: semantically similar tokens must be in the same partition, either green or red lists.
@@
    Cross-lingual semantic robust vocabulary partition: for semantically similar prefixes in different languages: “I watch” and “我 看”, the partitions of the vocabulary are the same.
Both Figure 5(b) and Figure 5(c) satisfy only one of the two factors, thus failing to recognize “电影” as a green token and losing cross-lingual consistency.
@@ 5.2 Defense Method against CWRA
We now improve the SIR so that it satisfies the two factors described above. As discussed in § 2.2, SIR uses the $\Delta$ function to represent vocabulary partition ($\Delta\in\mathbb{R}^{|\mathcal{V}|}$), where $\Delta_{i}>0$ indicating that $v_{i}$ is a green token. Based on eq. 4, SIR has already optimized for Factor 2 when using a multilingual embedding model. For prefixes $\boldsymbol{x}$ and $\boldsymbol{y}$, the similarity of their vocabulary partitions for the next token should be close to their semantic similarity:
@@
Based on SIR, we focus on Factor 1, i.e., cross-lingual semantic clustering of the vocabulary. Formally, we define semantic clustering as a partition $\mathcal{C}$ of the vocabulary $\mathcal{V}$: $\mathcal{C}=\{\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{|\mathcal{C}|}\}$ , where each cluster $\mathcal{C}_{i}$ consists of semantically equivalent tokens. Instead of assigning biases for each token in $\mathcal{V}$, we adapt the $\Delta$ function so that it yields biases to each cluster in $\mathcal{C}$, i.e., $\Delta\in\mathbb{R}^{|\mathcal{C}|}$. Thus, the process of adjusting the logits should be:
@@
where $C(i)$ indicates the index of $v_{i}$’s cluster within $\mathcal{C}$. By doing so, if token $v_{i}$ and $v_{j}$ are semantically equivalent, they will receive the same bias on logits:
$C(i)=C(j)\implies\Delta_{C(i)}=\Delta_{C(j)},$ (9)
i.e., if $v_{i}$ and $v_{j}$ are translations of each other, they will fall into the same list. To obtain such a semantic clustering $\mathcal{C}$, we treat each token in $\mathcal{V}$ as a node, and add an edge $(v_{i},v_{j})$ whenever $(v_{i},v_{j})$ corresponds to an entry in a bilingual dictionary. Therefore, $\mathcal{C}$ is all the connected components of this graph.
We name this method as X-SIR and evaluated it under the same setting as § 4. We also detail its limitations in § 8.
Method Text Summ. Question Ans.
[…]
SIR 0.431 66.18 0.321 71.42
X-SIR 0.554 43.49 0.507 34.98
Table 3: Cross-lingual consistencies in terms of PCC and RE under CWRA.
Figure 6: ROC curves of X-SIR and SIR.
Cross-lingual consistency & ROC curvers
Table 3 shows the cross-lingual consistency of SIR and X-SIR when confronted CWRA. X-SIR achieves significant improvements in terms of PPC and RE in both tasks. Consequently, as depicted in Figure 6, X-SIR substantially enhances the AUC under CWRA, with an increase in TPRs by ~0.4. Furthermore, X-SIR delivers performance on par with SIR in the absence of any attacks. These findings validate the two key factors of cross-lingual consistency that we identified in § 5.1.
@@
As shown in Table 4, X-SIR achieves better text quality than SIR in text summarization, and comparable performance on question answering, meaning the semantic partition of vocabulary will not negatively affect text quality.

### v2

@@ Abstract
Text watermarking technology aims to tag and identify content produced by large language models (LLMs) to prevent misuse. In this study, we introduce the concept of “cross-lingual consistency” in text watermarking, which assesses the ability of text watermarks to maintain their effectiveness after being translated into other languages. Preliminary empirical results from two LLMs and three watermarking methods reveal that current text watermarking technologies lack consistency when texts are translated into various languages. Based on this observation, we propose a Cross-lingual Watermark Removal Attack (CWRA) to bypass watermarking by first obtaining a response from an LLM in a pivot language, which is then translated into the target language. CWRA can effectively remove watermarks, decreasing the AUCs to a random-guessing level without performance loss. Furthermore, we analyze two key factors that contribute to the cross-lingual consistency in text watermarking and propose X-SIR as a defense method against CWRA.
@@ 1 Introduction
Large language models (LLMs) like GPT-4 OpenAI (2023) have demonstrated remarkable content generation capabilities, producing texts that are hard to distinguish from human-written ones. This progress has led to concerns regarding the misuse of LLMs, such as the risks of generating misleading information, impersonating individuals, and compromising academic integrity Chen and Shu (2023); Ai et al. (2024); Yuan et al. (2024); Xia et al. (2024). As a countermeasure, text watermarking technology for LLMs has been developed, aiming at tagging and identifying the content produced by LLMs Kirchenbauer et al. (2023a); Liu et al. (2023). Generally, a text watermarking algorithm embeds a message within LLM-generated content that is imperceptible to human readers, but can be detected algorithmically. By tracking and detecting text watermarks, it becomes possible to mitigate the abuse of LLMs by tracing the origin of texts and ascertaining their authenticity.
@@
The robustness of watermarking algorithms, i.e., the ability to detect watermarked text even after it has been modified, is important. Recent works have shown strong robustness under text rewriting and copy-paste attacks Liu et al. (2024b); Yang et al. (2023b). However, these watermarking techniques have been tested solely within monolingual contexts. In practical scenarios, watermarked texts might be translated He et al. (2024a, b, 2022); Jiao et al. (2023); Liang et al. (2023), raising questions about the efficacy of text watermarks across languages (see Figure 1). For example, a malicious user could use a watermarked LLM to produce fake news in English and then translate it into Chinese. Obviously, the deceptive impact persists regardless of the language, but it is uncertain whether the watermark would still be detectable after such a translation. To explore this question, we introduce the concept of cross-lingual consistency in text watermarking, aiming to characterize the ability of text watermarks to preserve their strength across languages. Our preliminary results on 2 LLMs $\times$ 3 watermarks reveal that current text watermarking technologies lack consistency across languages.
@@
In light of this finding, we propose the Cross-lingual Watermark Removal Attack (CWRA) to highlight the practical implications arising from deficient cross-lingual consistency. When performing CWRA, the attacker begins by translating the original language prompt into a pivot language, which is fed to the LLM to generate a response in the pivot language. Finally, the response is translated back into the original language. In this way, the attacker obtains the response in the original language and bypasses the watermark with the second translation step. CWRA outperforms re-writing attacks, such as re-translation and paraphrasing Liu et al. (2023), as it decreases the AUCs to a random-guessing level and achieves the highest text quality.
@@
To resist CWRA, we propose a defense method that improves the cross-lingual consistency of current LLM watermarking. Our method is based on two critical factors. The first is the cross-lingual semantic clustering of the vocabulary. Instead of treating each token in the vocabulary as the smallest unit when ironing watermarks, as done by KGW Kirchenbauer et al. (2023a), our method considers a cluster of tokens that share the same semantics across different languages as the smallest unit of processing. In this way, the post-translated token will still carry the watermark as it would fall in the same cluster as before translation. The second is cross-lingual semantic robust vocabulary partition. Inspired by Liu et al. (2024b), we ensure that the partition of the vocabulary are similar for semantically similar contexts in different languages. Despite its limitations, our approach (named X-SIR) substantially elevates the AUCs under the CWRA, paving the way for future research.
@@
  • Evaluation (§ 3): We reveal the deficiency of current text watermarking technologies in maintaining cross-lingual consistency.
  • Attack (§ 4): Based on this finding, we propose CWRA that successfully bypasses watermarks without degrading the text quality.
  • Defense (§ 5): We identify two key factors for improving cross-lingual consistency and propose X-SIR as a defense method against CWRA.
@@ 2.2 Watermarking for LMs
$s=(|\boldsymbol{x}|_{g}-\gamma|\boldsymbol{x}|)/\sqrt{|\boldsymbol{x}|\gamma(1-\gamma)},$ (2)
@@ 3.1 Definition
We define cross-lingual consistency as the ability of a text watermark to retain its strength after the text is translated into another language. We represent the original strength of the watermark as a random variable, denoted by $S$ (Appendix A), and its strength after translation as $\hat{S}$. To quantitatively assess this consistency, we employ the following two metrics.
@@
A lower RE indicates that the watermark retains strength close to its original value after translation, signifying great cross-lingual consistency.
@@ 3.2 Experimental Setup
We sampled a subset of 500 English prompts from the mc4 dataset Raffel et al. (2019) [^2] https://huggingface.co/datasets/mc4, and generated responses from the LLM using the text watermarking methods described in § 2.2. The default decoding method was multinomial sampling, and both the prompts and the LLM-generated responses were in English. To evaluate the cross-lingual consistency, these watermarked responses were translated into four languages using gpt-3.5-turbo-0613333https://platform.openai.com/docs/models: Chinese (Zh), Japanese (Ja), French (Fr), and German (De). Notably, English shares greater similarities with French and German, in contrast to its significant differences from Chinese and Japanese.
@@
  • Baichuan-7B Baichuan. (2023): an LLM trained on 1.2 trillion tokens. It offers bilingual support for both Chinese and English officially. We also found that its vocabulary covers Japanese tokens.
  • Llama-2-7B Touvron et al. (2023): trained on 2 trillion tokens and only provides support for English officially. Its vocabulary also contains tokens for European languages, such as German and French.
Method PCC $\uparrow$ RE (%) $\downarrow$
@@
Figure 2: Trends of watermark strengths with text length before and after translation. These are the average results of Baichuan-7B and Llama-2-7B. Given the distinct calculations for watermark strengths of the three methods, the y-axis scales vary accordingly.
@@ 3.3 Results
Table 1 presents the main results. We forced the model to generate 200 tokens in response following the setting of Kirchenbauer et al. (2023a). With response length restriction lifted, Figure 2 illustrates the trend of watermark strengths with text length.
@@
None of the three watermarking methods exhibits such a characteristic that its cross-lingual consistency between similar languages is significantly better than distant ones. This means that even if two languages have similar structures or shared words, it is still difficult for watermarks to transfer between them, which poses a big challenge to the cross-lingual consistency of watermarking.
@@
Overall, SIR indeed exhibits superior cross-lingual consistency compared to KGW and UW. It achieves the best average results across the two models and two metrics. When using Baichuan-7B, SIR notably outperforms other methods in terms of PCCs for all target languages. This finding highlights the importance of semantic invariance in preserving watermark strength across languages, which we will explore more in § 5. Despite its superiority, SIR still presents a notable reduction in watermark strength in cross-lingual scenarios, as evidenced by Figure 2c.
@@ 4.2 Results
CWRA demonstrates the most effective attack performance, significantly diminishing the AUCs and the TPRs. For one thing, existing watermarking techniques are not designed for cross-lingual contexts, leading to weak cross-lingual consistency. For another thing, strategies such as Re-translation and Paraphrase are essentially semantic-preserving text rewriting. Such strategies tend to preserve some n-grams from the original response, which may still be identifiable by the watermark detection algorithm. In contrast, CWRA reduces such n-grams due to language switching.
@@
Figure 4: ROC curves for KGW, UW, and SIR under various attack methods: Re-translation, Paraphrase and CWRA. We also present TPR values at a fixed FPR of 0.1. This is the overall result of text summarization and question answering. Figure 9(a) and Figure 9(b) display results for each task.
@@ 5.1 Two Key Factors of Cross-lingual Consistency
KGW-based watermarking methods fundamentally depend on the partition of the vocabulary, i.e., the red and green lists, as discussed in § 2.2. Therefore, cross-lingual consistency aims to achieve the following goal:
The green tokens in the watermarked text will still be recognized as green tokens after being translated into other languages.
With this goal in mind, we start our analysis with a toy case in Figure 5:
@@
[…]
@@
[…]
  3. 3.
[…]
Figure 5: Three cases of attempts to maintain cross-lingual consistency. Cross-lingual consistency can only be achieved if
[…]
is in the green list when watermark detection. Factor 1: semantically equivalent tokens should be in the same list (either red or green). In these cases,
[…]
[regular]
[…]
, and
[…]
[regular]
[…]
. Factor 2: the vocabulary partitions for semantically equivalent prefixes (
[…]
[regular]
and
[…]
) should be the same.
[…]
@@
    Cross-lingual semantic clustering of the vocabulary: semantically equivalent tokens must be in the same partition, either green or red lists.
@@
[…]
@@ 5.2 Defense Method against CWRA
We now improve the SIR so that it satisfies the two factors described above. In Figure 5, the tokens of the two languages are equivalent in one-to-one correspondence. In practice, we relax this semantic equivalence into semantic similarity.
As discussed in § 2.2, SIR uses the $\Delta$ function to represent vocabulary partition ($\Delta\in\mathbb{R}^{|\mathcal{V}|}$), where $\Delta_{i}>0$ indicating that $v_{i}$ is a green token. Based on eq. 4, SIR has already optimized for Factor 2 when using a multilingual embedding model. For prefixes $\boldsymbol{x}$ and $\boldsymbol{y}$, the similarity of their vocabulary partitions for the next token should be close to their semantic similarity:
@@
Based on SIR, we focus on Factor 1, i.e., cross-lingual semantic clustering of the vocabulary. Formally, we define semantic clustering as a partition $\mathcal{C}$ of the vocabulary $\mathcal{V}$: $\mathcal{C}=\{\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{|\mathcal{C}|}\}$ , where each cluster $\mathcal{C}_{k}$ consists of semantically similar tokens. Instead of assigning biases for each token in $\mathcal{V}$, we adapt the $\Delta$ function so that it yields biases to each cluster in $\mathcal{C}$, i.e., $\Delta\in\mathbb{R}^{|\mathcal{C}|}$. Thus, the process of adjusting the logits should be:
@@
where $C(i)$ indicates the index of $v_{i}$’s cluster within $\mathcal{C}$. By doing so, if token $v_{i}$ and $v_{j}$ are semantically similar, they will receive the same bias on logits:
$C(i)=C(j)\implies\Delta_{C(i)}=\Delta_{C(j)}.$ (9)
In other words, if $v_{i}$ and $v_{j}$ are translations of each other, they will fall into the same list. Algorithm 1 shows the procedure of semantic clustering. To obtain such a semantic clustering $\mathcal{C}$, we treat each token in $\mathcal{V}$ as a node, and add an edge $(v_{i},v_{j})$ whenever $(v_{i},v_{j})$ corresponds to an entry in a bilingual dictionary $\mathcal{D}$. Therefore, $\mathcal{C}$ is all the connected components of this graph.
Algorithm 1 Constructing semantic clusters
1:the vocabulary of the LM $\mathcal{V}$, a bilingual dictionary $\mathcal{D}$
2:Semantic clusters $\mathcal{C}$
3:Initialize a graph $G=(\mathcal{V},\varnothing)$
4:for each entry $(v_{i},v_{j})$ in $\mathcal{D}$ do
5:     if $v_{i}\in\mathcal{V}$ and $v_{j}\in\mathcal{V}$ then
6:         Add an edge $(v_{i},v_{j})$ to $G$
7:$\mathcal{C}=\varnothing$
8:for each connected component $\mathcal{C}_{k}$ in $G$ do
9:     $\mathcal{C}=\mathcal{C}\cup\{\mathcal{C}_{k}\}$
10:return $\mathcal{C}$
Setup
We name this method as X-SIR, implemented it with an unified external dictionary $\mathcal{D}$ that covers English (En), Chinese (Zh), Japanese (Ja), French (Fr), and German (De). Following the settings of § 3.2 and § 4.1, we apply X-SIR on Baichuan-7B, analyze it in the following and detail its limitations in § 8. We also present more results on other LLMs in Appendix C.
Cross-lingual consistency & ROC curves
Figure 6 shows the cross-lingual consistency of SIR and X-SIR when the original responses are translated into other languages. Considering both PCC and RE, X-SIR notably improves the cross-lingual consistency over SIR when the target language is Zh or Ja. However, for De and Fr, X-SIR improves only the PCC but not RE, which we regard as a limitation and discuss in § 8. Further, Figure 7 evaluates both methods’ watermark detection performance under cross-lingual scenarios. Whether directly translating the responses into other languages (Figure 7(a)) or using CWRA (Figure 7(b)), X-SIR substantially enhances the watermark detection performance, with an average increase in TPR by 0.448 and AUC by 0.221. These findings validate the two key factors of cross-lingual consistency that we identified in § 5.1.
Figure 6: Cross-lingual consistency in terms of PCC and RE. We follow the same setting as § 3.2 and plot 100%-RE for visualization.
(a) Direct translation
(b) CWRA
Figure 7: Watermark detection performance under cross-lingual scenarios. (a) Direct translation: the original English responses are translated into Ja or Zh following the setting in § 3.2. (b) CWRA: the full CWRA process as we did in § 4, where English is the original language and Chinese is the pivot language.
No attack & Paraphrase attack
The next question is whether the addition of semantic clustering of the vocabulary to X-SIR will affect the original detection performance of SIR. As shown in Figure 8, when there is no attack, both methods perform comparably. However, under the paraphrase attack, X-SIR performs slightly better than SIR. This is reasonable because paraphrasing can be regarded as a special kind of “translation” within the same language and X-SIR improves such intra-lingual consistency.
Figure 8: Watermark detection performance under no attack and paraphrase attack.
@@
As shown in Table 3, X-SIR achieves better text quality than SIR in text summarization, and comparable performance on question answering, meaning the semantic clustering of vocabulary will not negatively affect text quality.